OpenStudy (anonymous):

Find the value of three times the quantity x minus two when x equals eight. 1 eighteen twenty-two twenty-four

OpenStudy (oaktree):

That's\[3x-2\]Which is \[3(8) - 2 = 24-2 = 22\]

OpenStudy (anonymous):

thx